Please explain how to get this answer for me. Thanks:

A non volatile, non-ionic whose molecular weight is 140 is dissolved in the 875g of cyclohexane. The freezing point of this solution is 1.50 C. What mass of the compound was dissolved in the cyclohexane?
1
Expert's answer
2016-03-18T15:48:05-0400
Delta(T)=Cm*K (Delta(T)=T(freezing of pure solvent) - T(freezing of solution); Cm–Molality; K=20 – cryoscopic constant of cyclohexane)
Delta(T)=6.45-1.50=4.95
Cm=m/(M*m(solv))= x/(140*0.875) x-mass of the compound, M - molar mass of the compound, 0.875 - mass of the solvent in kg
x=Delta(T)*(M*m(solv))/K=30.31